# Quantum Entanglement of Anyonic Charges and Emergent Spacetime Geometry

### v1

@@ III Entanglement Structure via Mutual Information in a Disordered Potential
An instanton consisting of two fractional charges $e/2$, each residing on an opposite edge, gives rise to long-range entanglement. In this section, we investigate the spatial structure of this entanglement.
@@ III.1 Mutual Information and von Neumann Entropy
where $\lambda_{i}$ are the eigenvalues of the following correlation function (or equivalently, of the reduced density matrix of region DD[yang2021]
@@
Figure 3: (a) Mutual information between the two yellow boxes is plotted as a function of their position $x$ along the nanoribbon, measured from the left end. The size of each box is $(\ell,w)=(5,1)$, where $\ell$ is the length (number of sites along a zigzag edge) and $w$ is the width (number of carbon rows). (b) Mutual information as a function of the separation $\Delta y$ between two boxes, each with $(\ell,w)=(5,1)$. The maximal separation between two boxes is $\Delta y=W=16$. Results are presented for two scenarios: when each box contains a fractional charge, and when they do not. (c) Mutual information between two boxes, each with $(\ell,w)=(5,w)$, is plotted as a function of the width $w$. Here, the boxes are positioned to enclose a localized fractional charge at $x=120$. All three panels share the same parameters and disorder realization, with $(U,\Gamma)=(t,0.01t)$ and $(L,W)=(300,16)$. The horizontal dashed lines in lower row figures indicate zigzag edges.
In quantum information theory, mutual information quantifies the total correlations between two separate regions of a quantum system. It is defined in terms of entanglement entropy and is particularly well suited for probing non-local entanglement between spatially separated fractional charges, such as those localized on opposite zigzag edges. We compute the Hartree-Fock mutual information [mut_inf2023] between two regions $I$ and $J$, defined as
$M_{I,J}=S_{I}+S_{I}-S_{I,J}.$ (9)
Here, $S_{I}$, $S_{J}$, and $S_{I,J}$ denote the entanglement entropy of regions $I$, $J$, and their union $I\cup J$, respectively. We define the region II ($J$) as subregions. Mutual information depends not only on the distance between regions $I$ and $J$, but also on their sizes, reflecting the scale-dependent nature of entanglement.
@@ III.2 Mutual Information of an Instanton
We now investigate how non-local correlations are reflected in the mutual information. Figure 3(a) shows the mutual information between two boxes, each consisting of five sites along the zigzag edge, plotted as a function of their distance $x$ from the left end of the ribbon. Whenever a box contains a fractional charge $e/2$, corresponding to a site where the occupation number $\delta n_{m,\uparrow}$ changes abruptly (see Fig. 2), the mutual information exhibits a pronounced peak as a function of $x$. Because fractional charges are localized in random positions, the overall profile exhibits a sequence of alternating peaks and valleys.
In Fig. 3(b), we compare mutual information as a function of the vertical distance $\Delta y$ in two cases, depending on whether the horizontal box encloses a fractional charge or not. In both cases, each box consists of five sites along a single row, i.e. $\ell=5$ and $w=1$. In the first case, mutual information $M$ remains small until the regions begin to overlap with opposite zigzag edges at $\Delta y=W-1$, where a pronounced peak emerges. In contrast, in the second case, where the boxes do not contain fractional charges at the edges, no such peak is observed in $M$ as $\Delta y$ increases.
@@ III.3 Scale-Dependent Entanglement
We have also explored scale-dependent entanglement by computing mutual information for regions (boxes) that span multiple rows of carbon sites, as presented in Fig. 3(c). The results show clear signatures of non-local long-range correlations, as shown in Fig. 2. Together with Figs. 3(a) and 3(b), which present single-row configurations, these findings reveal the characteristic entanglement pattern surrounding a fractional charge $e/2$, schematically depicted in Fig. 4. This figure illustrates the scale-dependent entanglement structure associated with a fractionalized charge, where “scale” refers to the number of sites included in region $I$ or $J$. Note that short-range entanglement between $I$ and $J$ is excluded from $M_{I,J}$ because the two regions are spatially separated. The grouping of neighboring degrees of freedom into larger blocks allows one to probe entanglement at increasing length scales, a concept known as scale-dependent entanglement [Vidal2007]. The renormalization scale thereby defines a holographic direction along which bulk correlations decay exponentially [Swingle2012].
@@ IV Embedding of ZGNR in Hyperbolic-like Geometry
Although AdS spacetime does not directly emerge from the microscopic physics of condensed matter systems, a curved geometry (information geometry) with analogous features can arise through an embedding. A particularly relevant example is the soft-wall geometry [Andreev2006, Karch2006, Klebanov2008, Erlich2014], which introduces a smooth asymptotic boundary at $y=\pm W/2$ without imposing hard cutoffs, where $W$ is the ribbon’s width. Near these edges, the embedding generates a diverging metric that acts as an effective boundary, gradually suppressing the propagation of low-energy electrons toward $y=\pm W/2$. A non-isometric embedding captures the underlying information structure, such as patterns of correlation or entanglement, rather than the actual metric structure. (That is, it does not preserve distances or local geometry.)
@@
This construction reproduces the exponential stretching characteristic of the hyperbolic space, with the edges at $y\to\pm\infty$ effectively pushed to $y=\pm W/2$, realizing a soft-wall geometry in the emergent curved space as illustrated in Fig. 5(a). This defines an inhomogeneous spatial geometry that effectively models a warped AdS-like background. Here, $x$ runs along the length of the ribbon, while $y$ points into the bulk or embedding direction. The function $R(y)$ captures the exponential contraction characteristic of hyperbolic space as $|y|$ increases. It acts as a geometric shape function, describing how far the ribbon extends in the $x$-direction at a given depth $y$. Crucially, this is not a coordinate transformation, but an embedding function that defines the intrinsic geometry of the surface.
@@
$=\frac{W}{y_{0}\cosh^{2}(y/y_{0})}\ell.$
The geodesic distance along the ribbon decays exponentially as $y\to\infty$, reflecting the asymptotic narrowing of the geometry:
@@
This behavior is reminiscent of hyperbolic space, where distances contract or stretch exponentially with depth [weisstein2003poincare]. The corresponding Einstein field equations, which are provided in Appendix B, give a negative energy density, which is consistent with AdS.
Figure 5: (a) “Hyperbolic geometry” of a ZGNR with width $W=32$ and characteristic curvature scale $y_{0}=9$. A and B carbon sites are represented by red and blue colors, respectively. The ribbon comprises two interpenetrating triangular sublattices, shown in red and blue. The red and blue zigzag edge sites lie directly opposite each other across the ribbon. (b) The coarse-grained (disorder-averaged) mutual information $\log M(y)$ is computed with a transverse width $w=1$, for typical separations between fractional charges $\ell_{\text{frac}}=10,\,15$. The fitted slope yields an inverse decay length $\alpha=0.09$ (see Eq. (14)), with the curvature scale parameter fixed at $y_{0}=10$. Parameters for Hamiltonian are $(U,\Gamma)=(t,0.1t)$. The geometries of ribbon for $\ell_{\text{frac}}=10$ and $\ell_{\text{frac}}=15$ are $(L,W)=(200,48)$ and $(L,W)=(150,72)$, respectively. The results show little dependence on $\ell_{\text{frac}}$: the vertical separation between the two $\log M(y)$ curves corresponds to only a tiny relative difference, $\Delta M(y)\sim 10^{-3}$.
@@ V Emergent Geodesic Distance from Coarse-Grained Mutual Information
Disorder induces pronounced microscopic inhomogeneities in mutual information along the $x$ axis, as shown in Fig. 3(a). Fractional charges appear irregularly along the zigzag edges, and their number scales linearly with ribbon length. (Conformal invariance, dependent on scale and translational symmetry, is explicitly broken.) To probe the emergent space at larger scales, we coarse-grain the mutual information along the $x$-axis over a length scale several times larger than the average horizontal spacing between neighboring fractional charges. Averaging over many disorder realizations serves as an effective coarse-graining procedure, since the positions of fractional charges along the edges vary randomly in each realization. The resulting averaged mutual information is independent of $x$ but retains a dependence on $y$ (see the definition of $y$ below Eq. (11)).
In our embedded space, the coordinate $y$ maps to the radial depth in the bulk: small $y$ corresponds to deeper layers (infrared), whereas large $|y|$ indicates proximity to the boundary. Although this mapping inverts the usual AdS convention, it meaningfully reflects the physical structure of zigzag nanoribbons, where fractional charges are edge-localized. Each pair of mirrors at $y$ and y-y is treated as lying at the same bulk depth $y$; the emergent geometry depends only on the radial distance, not on the sign of $y$. In hyperbolic space, the depth is scalar: there is no distinction between $\pm y$ in terms of curvature or geodesic distance.
We have computed the mutual information averaged with disorder $M(y)$ as a function of the hyperbolic distance between regions $I$ and $J$, see Fig. 5(b). The resulting mutual information $M(y)$ is well described a by the following expression
@@
where the effective hyperbolic distance $d_{\mathrm{geo}}(y)$ is given by Eq. (12) with $\ell$ being replaced by $\ell_{\text{frac}}$, which characterizes the typical spatial distance between fractional charges, and $\alpha$ is a certain positive constant. This coarse-grained mutual information defines a geodesic distance in the information geometry - an emergent feature shaped by correlations rather than by explicit spatial embedding. The coarse-grained mutual information $M(y)$ captures the nonlocal entanglement between fractional charges. The exponential decay along the $y$-axis mirrors the behavior of correlation functions in holographic settings such as AdS/CFT, where geodesic lengths or minimal surfaces determine the strength of correlations [Ryu2006] (see Appendix C). This justifies interpreting the coarse-grained, disorder-averaged system as realizing an emergent hyperbolic-like geometry.
Our results suggest that the emergent space reconstructed from the entanglement correlations behaves as if it is hyperbolic. This perspective aligns with Swingle’s proposal [Swingle2012] and the broader entanglement = geometry program initiated by Van Raamsdonk [van2010]. As in AdS/CFT, the decay of correlations with increasing geodesic length reflects an underlying emergent geometry. Here, the bulk geometry arises not from fundamental curvature but from the entanglement structure itself, revealing a spatial slice of a higher-dimensional curved spacetime emergent from quantum information.
@@ VI Summary
In this work, we examined the highly entangled (strongly correlated) ground state of ZGNRs and shown that their behavior can be effectively described in terms of an emergent space-time geometry. In particular, we identified the regimes and conditions under which this emergent geometry remains robust, providing criteria for its stability and persistence. A key diagnostic is the mutual information between anyon fractional charges located on opposite zigzag edges, which serves as a sensitive probe of the underlying entanglement structure. Its spatial profile encodes an effective bulk metric $g_{\mu\nu}(x,y)$, thereby providing a coarse-grained mapping from the underlying microscopic lattice to an emergent geometry reminiscent of hyperbolic space. In this description, edge-to-edge mutual information plays the role of a geodesic distance in a negatively curved space, reminiscent of the AdS/CFT correspondence. The scale-dependent decay of mutual information thus substantiates the emergence of an AdS-like bulk geometry, offering a concrete realization of holography in a system that lacks conformal symmetry or intrinsic curvature, yet still manifests the deep connection between entanglement and geometry. Data supporting the findings of this article are openly available [le_2025_16916229].

### v2

@@ III Entanglement Structure via Mutual Information in a Disordered Potential
An instanton consisting of two fractional charges $e/2$, each residing on an opposite edge, gives rise to long-range entanglement. In this section, we investigate the spatial structure of this entanglement via mutual information.
@@ III.1 Mutual Information and von Neumann Entropy
where $\lambda_{i}$ are the eigenvalues of the following correlation function (or equivalently, of the reduced density matrix of region DD[26]
@@
The entanglement entropy of a spatial region $D$ contains a universal subleading constant term $\gamma$, known as the TEE [39, 29]:
$S_{D}=\alpha L-\gamma,$ (9)
where $L$ denotes the length of the boundary of the region $D$. The coefficient $\gamma$ is a universal constant, independent of microscopic details, and characterizes the topological order in a certain region of the parameter space. The TEE is a topological invariant in the graphene ribbon system. Disorder must be present both at the edges and in the bulk; otherwise, one does not obtain a universal value of the TEE that is independent of the shape of any smooth domain within the ribbon. This is a stringent requirement. The region $D$ used to compute the TEE must be chosen sufficiently far from the edges in order to suppress non-universal edge contributions and reliably extract the topological term [26].
Defining the von Neumann entanglement entropy for identical fermions is generally subtle, due to the non-factorizable structure of the many-body Fock space and the resulting ambiguity in partial traces (see, e.g., Ref. [3]). In the present work, however, this issue does not arise because we consider spatial bipartitions on a lattice, for which the fermionic Fock space factorizes into a tensor product of local mode algebras. The ambiguity occurs only for particle-based partitions, whereas for spatial bipartitions the reduced density matrix and entanglement entropy are well defined [60, 58].
In quantum information theory, mutual information quantifies the total correlations between two separate regions of a quantum system. It is defined in terms of entanglement entropy and is particularly well suited for probing non-local entanglement between spatially separated fractional charges, such as those localized on opposite zigzag edges. We compute the Hartree-Fock mutual information [37] between two regions $I$ and $J$, defined as
$M_{I,J}=S_{I}+S_{J}-S_{I\cup J}.$ (10)
Here, $S_{I}$, $S_{J}$, and $S_{I\cup J}$ denote the entanglement entropy of regions $I$, $J$, and their union $I\cup J$, respectively. We define the region II ($J$) as sub-regions. Mutual information depends not only on the distance between the regions $I$ and $J$, but also on their sizes, reflecting the scale-dependent nature of entanglement. The definition of the mutual information is valid for both SPT phases and intrinsically topologically ordered phases, independent of whether disorder is present.
@@ III.2 Mutual Information of an Instanton
We now investigate how non-local correlations are reflected in the mutual information. Figure 3(a) shows, in the weak disorder regime, the mutual information between two boxes, each consisting of five sites along the zigzag edge, plotted as a function of their distance $x$ from the left end of the ribbon. The box size is chosen to optimally reveal the behavior of $M$ as a function of $x$ and $y$, while reflecting the fact that semions are localized on the zigzag edges, with their probability density decaying into the bulk within a few lattice constants. The horizontal and vertical dimensions of each box are chosen to be comparable to the size of the semions, which sets the relevant length scales of the problem. Whenever a box contains a fractional charge $e/2$, corresponding to a site where the occupation number $\delta n_{m,\uparrow}$ changes abruptly (see Fig. 2(b)), the mutual information exhibits a pronounced peak as a function of $x$. Because fractional charges are localized in random positions, the overall profile exhibits a sequence of alternating peaks and valleys.
Figure 3: (a) Mutual information between two yellow boxes as a function of their position $x$, measured from the left end (see inset), obtained from a single disorder realization. Each box has size $(\ell,w)=(5,1)$, where $\ell$ is the length along a zigzag edge and $w$ is the width measured in carbon rows. Here $(U,\Gamma)=(t,0.01t)$. The red horizontal line indicates $M(x)$ of the translationally invariant, disorder-free case $(\Gamma=0)$. Horizontal dashed lines in the insets indicate zigzag edges. (b) Mutual information at $x=120$ (hosting a fractional charge) and $x=71$ (without a fractional charge), plotted as a function of the vertical separation $\Delta y$ between the two boxes, with a maximum separation of $\Delta y=W=16$. Here $(U,\Gamma)=(t,0.01t)$. For comparison, we also show the mutual information of the disorder-free system with $(U,\Gamma)=(t,0)$. (c) Mutual information as a function of box width $w$, with boxes positioned to enclose a localized fractional charge at $x=120$. (d) Mutual information versus position $x$ in the strong-disorder regime $(U,\Gamma)=(t,12t)$, obtained from a single disorder realization. (e) Mutual information plotted as a function of $\Delta y$ at two selected positions in panel (d). These curves are representative of the typical behavior at various values of $x$. (f) Schematic illustration of the probability density of edge states with and without disorder. In the absence of disorder, the single-particle state is localized on either zigzag edge at $y=\pm W/2$ (perpendicular to the zigzag edge direction), while remaining delocalized along the zigzag edge. In the presence of weak disorder, the single-particle edge state becomes fractionalized into two spatially separated halves, each localized on one of the two edges. This state is localized both perpendicular to and along the zigzag direction. All numerical calculations are performed for a ribbon with ($L,W)=(300,16)$.
Figure 3(b) shows examples of $M(\Delta y)$ vs $\Delta y$ at some selected values of $x$. They compare mutual information as a function of the vertical distance $\Delta y$ in two cases, depending on whether the horizontal box encloses a fractional charge or not. In both cases, each box consists of five sites along a single row, i.e. $\ell=5$ and $w=1$. In the first case, the mutual information $M$ remains small up to $\Delta y=W-1$, where a pronounced peak emerges. In contrast, in the second case, where the boxes do not contain fractional charges at the edges, no such peak is observed in $M$ as $\Delta y$ increases. Figure 3(b) also shows the disorder-free case, in which no peak is found either. In other words, in the absence of semions, no peak appears at $\Delta y=W-1$ and the edge wave functions do not spatially split between opposite edges (see Fig. 3(f)). Figure 3(e) displays the mutual information $M(\Delta y)$ at two values of $x$ in the strong-disorder regime. They do not exhibit a pronounced peak at the edges. The absence of semions in the strong-disorder regime is consistent with the observation that the gap in the density of states closes and the topological order is destroyed [34].
@@ III.3 Scale-Dependent Entanglement
We have also explored scale-dependent entanglement by computing mutual information for regions (boxes) that span multiple rows of carbon sites, as presented in Fig. 3(c). Together with Figs. 3(a) and 3(b), which present single-row configurations, these findings reveal the characteristic entanglement pattern surrounding a fractional charge $e/2$, schematically depicted in Fig. 4. This figure illustrates the scale-dependent entanglement structure associated with a fractionalized charge, where “scale” refers to the number of sites included in region $I$ or $J$. Note that short-range entanglement between $I$ and $J$ is excluded from the mutual information $M_{I,J}$, since the two regions are spatially separated on opposite sides of the system. The grouping of neighboring degrees of freedom into larger blocks allows one to probe entanglement at increasing length scales, a concept known as scale-dependent entanglement [55]. The renormalization scale thereby defines a holographic direction along which bulk correlations decay exponentially [52].
@@ IV Embedding of ZGNR in Hyperbolic-like Geometry
Although AdS spacetime does not directly emerge from the microscopic physics of condensed matter systems, a curved geometry (information geometry) with analogous features can arise through an embedding. A particularly relevant example is the soft-wall geometry [2, 25, 30, 12], which introduces a smooth asymptotic boundary at $y=\pm W/2$ without imposing hard cutoffs, where $W$ is the ribbon’s width. Near these edges, the embedding generates a diverging metric that acts as an effective boundary, gradually suppressing the propagation of low-energy electrons toward $y=\pm W/2$. A non-isometric embedding captures the underlying information structure, such as patterns of correlation or entanglement, rather than the actual spacetime metric structure. (That is, it does not preserve distances or local geometry.)
@@
This construction reproduces the exponential stretching characteristic of the hyperbolic space, with the edges at $y\to\pm\infty$ effectively pushed to $y=\pm W/2$, realizing a soft-wall geometry in the emergent curved space as illustrated in Fig. 5(a). This defines an inhomogeneous spatial geometry that effectively models a warped AdS-like background. Recall that $x$ runs along the length of the ribbon, while $y$ points into the bulk or embedding direction. The function $R(y)$ captures the exponential contraction characteristic of hyperbolic space as $|y|$ increases. It acts as a geometric shape function, describing how far the ribbon extends in the $x$-direction at a given depth $y$. Crucially, this is not a coordinate transformation, but an embedding function that defines the intrinsic geometry of the surface.
@@
$=\frac{W}{y_{0}\cosh^{2}(y/y_{0})}\ell.$ (14)
(We take $y=0$ to be the center of the ribbon.) The geodesic distance along the ribbon decays exponentially as $y\to\infty$, reflecting the asymptotic narrowing of the geometry:
@@
This behavior is reminiscent of hyperbolic space, where distances contract or stretch exponentially with depth [56]. The corresponding Einstein field equations, which are provided in Appendix B, give a negative energy density, which is consistent with AdS geometry.
Figure 5: (a) “Hyperbolic geometry” of a ZGNR with width $W=32$ and characteristic curvature scale $y_{0}=9$. A and B carbon sites are represented by red and blue colors, respectively. The ribbon comprises two interpenetrating triangular sublattices, shown in red and blue. The red and blue zigzag edge sites lie directly opposite each other across the ribbon. (b) The coarse-grained (disorder-averaged) mutual information $\log M(y)$ is computed with a transverse width $w=1$, for typical separations between fractional charges $\ell_{\text{frac}}=10,\,15$. Larger $d_{\mathrm{geo}}(y)$ corresponds to points closer to the ribbon center. The fitted slope (dashed lines) yields an inverse decay length $\alpha=0.09$, with prefactors $A=0.006$ for the red data points and $A=0.0035$ for the blue data points (see Eq. (16)), with the curvature scale parameter fixed at $y_{0}=10$. The blue and red data points were computed with parameters in topological order regime $(U,\Gamma)=(t,0.1t)$, while the black points correspond to strong-disorder regime $(U,\Gamma)=(t,12t)$. The geometries of ribbon for $\ell_{\text{frac}}=10$ and $\ell_{\text{frac}}=15$ are $(L,W)=(200,48)$ and $(L,W)=(150,72)$, respectively. The results in topological order regime show little dependence on $\ell_{\text{frac}}$: the vertical separation between the two $\log M(y)$ curves corresponds to only a tiny relative difference, $\Delta M(y)\sim 10^{-3}$. The dot-dashed line represent $M(y)$ in the strongly disordered regime.The error bars are estimated from the variance of the mutual information computed over 89 (blue and red data points) and 50 (black data points) disorder realizations.
@@ V Emergent Geodesic Distance from Coarse-Grained Mutual Information
Disorder induces pronounced microscopic inhomogeneities in mutual information along the $x$ axis, as shown in Fig. 3(a). Fractional charges appear irregularly along the zigzag edges, and their number scales linearly with ribbon length. (Conformal invariance, dependent on scale and translational symmetry, is explicitly broken.) To probe the emergent geometry at larger scales, we coarse-grain the mutual information along the $x$-axis over a length scale several times larger than the average horizontal spacing between neighboring fractional charges. Averaging over many disorder realizations serves as an effective coarse-graining procedure, since the positions of fractional charges along the edges vary randomly in each realization. The resulting averaged mutual information is independent of $x$ but retains a dependence on $y$ (see the definition of $y$ below Eq. (12)).
In our embedded space, the coordinate $y$ maps to the radial depth in the bulk: small $y$ corresponds to deeper layers (infrared), whereas large $|y|$ indicates proximity to the boundary. Although this mapping inverts the usual AdS coordinate convention, it meaningfully reflects the physical structure of zigzag nanoribbons, where fractional charges are edge-localized. Each pair of mirrors at $y$ and y-y is treated as lying at the same bulk depth $y$; the emergent geometry depends only on the radial distance, not on the sign of $y$. In hyperbolic space, the depth is scalar: there is no distinction between $\pm y$ in terms of curvature or geodesic distance.
We have computed the averaged mutual information $M(y)$ over many disorder realizations and over horizontal coordinates $x$, which results in a function of the hyperbolic distance between regions $I$ and $J$, see Fig. 5(b). It is well described by the following expression:
@@
where the effective hyperbolic distance $d_{\mathrm{geo}}(y)$ is given by Eq. (14) with $\ell$ being replaced by $\ell_{\text{frac}}$, which characterizes the typical spatial distance between fractional charges, and $\alpha$ is a certain positive constant (see the caption of Fig. 5(b) for a particular value of $\alpha$ and $A$). This coarse-grained mutual information defines a geodesic distance in the information geometry - an emergent feature shaped by correlations rather than by explicit spatial embedding. The coarse-grained mutual information $M(y)$ captures the nonlocal entanglement between fractional charges. The exponential decay along the $y$-axis mirrors the behavior of correlation functions in holographic settings such as AdS/CFT, where geodesic lengths or minimal surfaces determine the strength of correlations [46] (see Appendix C). This justifies interpreting the coarse-grained, disorder-averaged system as realizing an emergent hyperbolic-like geometry. Note that, in the absence of semions or in the strong disorder regime, the geometry is qualitatively different. In particular, the AdS geometry does not emerge, since the exponential behavior fails to develop; see Fig. 5(b). In the parameter regime where the topological entanglement entropy is universal, we find that the constant $\alpha$ is also universal. Within numerical accuracy, it exhibits no dependence on the disorder strength $\Gamma$ or the on-site interaction $U$.
Our results suggest that the emergent space reconstructed from the entanglement correlations behaves as if it is hyperbolic. This perspective aligns with Swingle’s proposal [52] and the broader entanglement = geometry program initiated by Van Raamsdonk [53]. As in AdS/CFT, the decay of correlations with increasing geodesic length reflects an underlying emergent geometry. Here, the bulk geometry arises not from fundamental curvature but from the entanglement structure itself, revealing a spatial slice of a higher-dimensional curved spacetime emergent from quantum information.
@@ VI Summary
In this work, we examined the highly entangled (strongly correlated) ground state of ZGNRs and shown that their behavior can be effectively described in terms of an emergent space-time geometry. In particular, we identified the regimes and conditions under which this emergent geometry remains robust, providing criteria for its stability and persistence. A key diagnostic is the mutual information between anyon fractional charges located on opposite zigzag edges, which serves as a sensitive probe of the underlying entanglement structure. Its spatial profile encodes an effective bulk metric $g_{\mu\nu}(x,y)$, thereby providing a coarse-grained mapping from the underlying microscopic lattice to an emergent geometry reminiscent of hyperbolic space. In this description, edge-to-edge mutual information plays the role of a geodesic distance in a negatively curved space, reminiscent of the AdS/CFT correspondence. The scale-dependent decay of mutual information thus substantiates the emergence of an AdS-like bulk geometry, offering a concrete realization of holography in a system that lacks conformal symmetry or intrinsic curvature, yet still manifests the deep connection between entanglement and geometry.
Anyonic phenomena in quasi-one-dimensional systems provide an instructive contrast to those found in genuinely two-dimensional topological phases. In particular, two-leg electron ladders hosting Abelian anyons belong to the same universality class as zigzag graphene nanoribbons [9] and share a character that cannot be classified within the framework of modular tensor categories (MTCs) [28], in sharp contrast to intrinsically topologically ordered two-dimensional systems. In (2+1)-dimensional gapped phases with intrinsic topological order, quasiparticle excitations obey fractional statistics and possess fusion and braiding properties that satisfy the axioms of a modular tensor category; hence their excitations are naturally described by an MTC. Related anyon-inspired properties have also been explored in other one-dimensional platforms, including Rydberg atom lattices [38] and supersymmetric lattice models [23], where the Hilbert space is constrained to realize the fusion algebra of non-Abelian anyons. While these systems do not support genuine anyonic braiding or intrinsic topological order, they nevertheless exhibit nontrivial entanglement structures, highlighting the richness of anyon-inspired physics beyond strictly two-dimensional settings.
Traditionally, anyons have been classified primarily through their quantum statistical phases, namely their exchange properties. In contrast, our approach examines semions from an information-theoretic perspective, focusing on the structure of the hidden “spacetime” encoded in their entanglement patterns. This viewpoint naturally raises several intriguing questions. Can anyons be systematically classified according to the information-theoretic geometries they generate? Do different types of anyons give rise to distinct emergent geometrical structures? Disordered zigzag graphene nanoribbons exhibit Abelian semionic structure already at the level of the ground state. In contrast, in physical realizations described by MTCs, anyons, potentially including non-Abelian species, arise as quasiparticle excitations above the ground state. Since the computation of entanglement for excited states and for systems supportings non-Abelian anyons is subtle and lacks a general framework [3], an important open question is how features typically associated with MTC descriptions are encoded in the emergent entanglement geometry. Addressing this question across different intrinsically topologically ordered phases may provide a new information-theoretic route to understanding topological phases of matter.
Data supporting the findings of this article are openly available [33]. Numerical calculations were performed using HFGN code [35].